Session: Nov 2010 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Franco Fiocchi at Mon Nov 01 15:06:19 2010) Success:

Hartebeesthoek (Jonathan Quick at Mon Nov 01 15:19:31 2010) Success:
No problems noted

Torun (K. Borkowski at Tue Nov 02 07:22:22 2010) Success:
There were temporary problems with one of BBCs.

Effelsberg (Uwe Bach at Tue Nov 02 08:52:28 2010) Failure:RECE
A broken switch in the receiver had to be replaced. Was done during the NME, but finished too late to participate.

Onsala (Michael Lindqvist at Tue Nov 02 14:28:47 2010) Minor failures: RFI
Severe RFI. Our new broad band IF-system seems to be much more sensitive to RFI at L-band than anticipated.

Westerbork (Geert Kuper at Tue Nov 02 17:56:59 2010) Success:
No known problems